# FREE-Edit: Using Editing-aware Injection in Rectified Flow Models for Zero-shot Image-Driven Video Editing

### 01 Mar 2026 v1

@@ Abstract
Image-driven video editing aims to propagate edit contents from the modified first frame to the rest frames. The existing methods usually invert the source video to noise using a pre-trained image-to-video (I2V) model and then guide the sampling process using the edited first frame. Generally, a popular choice for maintaining motion and layout from the source video is intervening in the denoising process by injecting attention during reconstruction. However, such injection often leads to unsatisfactory results, where excessive injection leads to conflicting semantics from the source video while insufficient injection brings limited source representation. Recognizing this, we propose an Editing-awaRE (REE) injection method to modulate injection intensity of each token. Specifically, we first compute the pixel difference between the source and edited first frame to form a corresponding editing mask. Next, we track the editing area throughout the entire video by using optical flow to warp the first-frame mask. Then, editing-aware feature injection intensity for each token is generated accordingly, where injection is not conducted on editing areas. Building upon REE injection, we further propose a zero-shot image-driven video editing framework with recent-emerging rectified-Flow models, dubbed FREE-Edit. Without fine-tuning or training, our FREE-Edit demonstrates effectiveness in various image-driven video editing scenarios, showing its capability to produce higher-quality outputs compared with existing techniques. Project page: https://free-edit.github.io/page/.
Figure 1: We propose an Editing-awaRE (REE) feature injection method for pre-trained rectified Flow models to conduct zero-shot image-driven video editing, dubbed FREE-Edit. Given the edited first frame, our method generates the output video, which propagates editing contents while leaving other parts unchanged (e.g., motion). Here, the commonly employed vanilla feature injection always leads to conflicting semantics with the edited first frame. We also provide the w/o injection results, which cannot preserve motion well. In contrast, our REE feature injection method modulates the injection intensity of each token adaptively, which achieves the most elegant results.
@@ 1 Introduction
Given the above considerations, we propose an Editing-awaRE (REE) injection method, which modulates the injection intensity of each token according to whether it is edited or not. Specifically, we design a modulation weight $\bm{\lambda}$ to implement a linear interpolation with Query $\mathbf{Q}$} (or Key $\mathbf{K}$}) in the self-attention of the reconstruction branch and the corresponding $\tilde{{\mathbf{Q}}}$ (or $\tilde{{\mathbf{K}}}$) of the editing branch. To compute the modulation weight $\bm{\lambda}$, we first use differences of the source and edited first frame to mask the editing area. Next, we track the editing area throughout the entire video by evaluating optical flow [27] for the source video, which is used to warp the first-frame mask. Then, the modulation weight $\bm{\lambda}$ can be obtained by inversing the warped mask sequence, where no injection is performed in the editing area to maintain appearance modification.
We apply the proposed REE injection method into the pre-trained Flow model LTX-Video [8] to form a training-free image-driven video editing framework, which inspires the name of our method: FREE-Edit. Though simple, the extensive evaluation of FREE-Edit demonstrates our capability of generating high-quality appearance- and temporally consistent output. For example, as seen in Fig. 1, motion cannot be preserved well in FREE-Edit w/o-injection due to the limited motion signal. Besides, FREE-Edit w/ vanilla feature injection always leads to conflicting semantics from the source video with the edited appearance. In contrast, the proposed REE injection yields the most visually consistent results. To sum up, our contributions are:
@@ 2 Related Works
Text-driven Video Generation and Editing. As for Text-to-Video (T2V),  [1, 3] insert temporal layers a pretrained Text-to-Image (T2I) model [24] to obtain temporal modeling. Another line of methods [8, 32] employs 3D VAE to incorporate 3D convolutions, which operate videos both spatially and temporally. Besides, text-driven video editing expects to generate an edited video that adheres to the target prompt while the original motion and semantic layout are preserved. Many efforts [31, 22, 7, 14, 16] expand the U-Net architecture of a pre-trained T2I diffusion model (e.g., SD [24]) along the temporal axis for 3D correspondence, where the denoising process is performed under the target-prompt guidance. Recent progress [29] has also included RF models to perform editing.
Image-to-Video Generation and Editing. Several representative works [2, 8, 32] establish the foundation for image-to-video (I2V). Then, a group of methods [8, 28] propose a first-frame conditioning model with rectified-flow training procedure. Besides, some techniques [12, 6, 19] use these I2V models to carry out image-driven video editing by propagating the first edit frame. Here, besides training-free AnyV2V [12] and VideoShop [6], I2VEdit [19] performes finetuning on temporal attention layers for coarse motion extraction, which however, is time- and resource-consuming. Moreover, very recent work Go-with-the-Flow [4] proposes a motion-controlled method by fine-tuning video diffusion models, which also allows for first-frame guided video editing.
@@
Figure 2: The pipeline illustration of our FREE-Edit. Top: It obeys an “inversion-then-editing” pipeline. Starting from inverted noisy latent ${\bm{z}}_{1}$, the reconstructed video ${\bm{z}}_{0}$ and edited video $\tilde{{\bm{z}}}_{0}$ takes the source ${{\mathbf{X}}}^{1}$ and edited first frame $\hat{{\mathbf{X}}}^{1}$ as condition signal, respectively. Bottom: We design an Editing-awaRE (REE) injection method, which designs a modulation weight $\bm{\lambda}$ to adaptively replaces the intermediate model representations ($\tilde{{\mathbf{Q}}}$ and $\tilde{{\mathbf{K}}}$) in the editing process with those ($\mathbf{Q}$} and $\mathbf{K}$}) in the reconstruction process through self-attention blocks. Here, we first use optical flow to warp the automatically calculated first-frame editing mask, which yields tracked editing masks for subsequent frames. Based on them, we compute the modulation weight $\bm{\lambda}$ for each token, where no injection is performed in the editing area.
@@ 3 Preliminary
After the velocity field $v_{\theta}$ is trained, pure noise ${\bm{z}}_{1}$ is gradually recovered as a clean latent ${\bm{z}}_{0}$, where the sampling process can be described using a series of $N$ discrete timesteps $t_{i}$:
${\bm{z}}_{t_{i-1}}={\bm{z}}_{t_{i}}+(t_{i-1}-t_{i})v_{\theta}({\bm{z}}_{t_{i}},t_{i}),$ (3)
Transformer Model for Image-to-Video. LTX-Video [8] is an open-source real-time generative model for I2V and T2V. Specifically, it uses a training procedure based on RF with 28 DiT transformer blocks [20]. Besides, it employs 3D causal VAEs to deal with spatio-temporal representation, where a $(1+rn)$-frame video is compressed into $(1+n)$ latent frames with temporal compression ratio $r$. Here, the first frame is encoded as a separate latent frame. Besides, LTX-Video can perform real-time generation due to its highly compressed latent space.
@@ 4.1 Vanilla Feature Injection on RF Models
Given that the motion signal from the inverted noise alone is limited, AnyV2V[12] introduces a vanilla feature injection method in the “inversion-then-editing” pipeline. Next, we will present how to reproduce it to match RF video models. Formally, we use $\mathbf{A}$} and $\tilde{{\mathbf{A}}}$ to denote the full spatiotemporal self-attention maps in source denoising (reconstruction) and editing denoising trajectory, respectively. Then, to obtain more layout and motion representation from the source video, vanilla injection applies the attention maps $\mathbf{A}$}} to replace attention maps $\tilde{{\mathbf{A}}}$, where $\tilde{{\mathbf{A}}}$ and $\mathbf{A}$} are obtained by:
@@ 4.2 The Proposed REE Injection and FREE-Edit
The remaining problem is how to use the editing mask sequence ${\mathbf{M}}=\{{\mathbf{M}}^{0},{\mathbf{M}}^{1},...,{\mathbf{M}}^{rn}\}$ to generate modulation weight $\bm{\lambda}$. Since the temporal compression of the RF I2V model, we take the mask of the $\left\lceil\frac{r}{2}\right\rceil$-th frame for each $r$-frame video chunk. Thus, the compressed editing mask sequence ${\mathbf{M}}_{\mathbf{c}}=\{{\mathbf{M}}^{0},{\mathbf{M}}^{\left\lceil\frac{r}{2}\right\rceil},...,{\mathbf{M}}^{\left\lceil\frac{r}{2}\right\rceil\cdot n}\}$ is with $(1+n)$ mask frames. Then, to match the dimension of Query and Key in Eq. (7), we further down-sample and flatten each mask frame in ${\mathbf{M}}_{\mathbf{c}}$ into size of $l\times 1$, yielding ${\mathbf{M}}_{c\cap d}\in\mathbb{R}^{(1+n)\times l\times 1}$. Finally, our modulation weight $\bm{\lambda}$ can be obtained by inversing the ${\mathbf{M}}_{c\cap d}$:
@@
FREE-Edit with the proposed REE Injection. In Fig. 2, we propose a training-free image-driven video editing framework, dubbed FREE-Edit, which obeys an “inversion-then-editing” pipeline. The core is our REE injection, which adaptively replaces $\tilde{{\mathbf{Q}}}$ and $\tilde{{\mathbf{K}}}$ in the editing trajectory with $\mathbf{Q}$}} and $\mathbf{K}$}} in the reconstruction editing trajectory through self-attention blocks in transformers. Here, clean tokens (i.e., source reference tokens and edited reference tokens corresponding to ${{\mathbf{X}}}^{1}$ and $\hat{{\mathbf{X}}}^{1}$) are combined with noisy tokens in each time step.
Note that previous text-driven technique Fatezero [22] also explores attention blending with masks, which are generated from cross-attention maps of the text token associated with the foreground object. However, cross-attention maps struggle to indicate editing regions in certain editing types (e.g., adding objects), where the location of the added object is uncertain. In contrast, we use optical flow to track editing regions, which can be well adapted to various editing types.
Besides, our method differs from TokenFlow [7], which propagates diffusion features during editing based on inter-frame correspondences calculated during inversion. Specifically, TokenFlow performs correspondence propagation in text-driven editing, whereas FREE-Edit employs editing-mask propagation in image-driven editing.
@@ 5.1 Experimental Setup
Implementation Details. We conduct all experiments on a single Tesla A100 GPU. Besides, we use the classifier-free guidance (CFG) [10] of 3 during the editing process, and that of 1 in the inversion process, which is under the same negative prompt “worst quality, inconsistent motion, blurry, jittery, distorted”. Additionally, we adopt 50 time steps as noise schedule. We set $thr=35$ in Eq. (9) empirically. We use LTX-Video-2B [8] for the base I2V model for its real-time performance. As for feature injection, we conduct Eq. (7) in all transformer blocks and all time steps.
I2V-Edit-Benchmark. Considering there is no publicly available dataset for image-driven video editing, this paper constructs a mini benchmark named I2V-Edit-Bench to evaluate the proposed method. Specifically, it contains 60 videos and their corresponding edited first frames. The videos are collected from the Internet website, and DAVIS dataset [21], and their edit frames are created by using professional software (e.g., Photoshop) and MagicBrush editing tool [33]. Each video consists of 41 to 200 frames. We provide 6 editing types: including object replacement, style transfer, adding or removing objects, changing attributes, or background editing.
@@
Evaluation Metrics of Edited Videos. We use CLIP score ($\uparrow$) to measure editing fidelity, which calculates the similarity of CLIP image embedding [23] of each edited frame and the given first edit frame. Besides, we apply Warp Error ($\downarrow$[7, 14] to reflect temporal consistency. Specifically, we estimate optical flow [27] of the input video to warp the edited frames and compute average MSE between warped frames and the target ones. Moreover, with our generated mask sequence $\mathbf{M}$}, we measure the fidelity of the non-editing area by five metrics: CLIP score+ ($\uparrow$), Warp Error+ ($\downarrow$), SSIM+ [30] ($\uparrow$), PSNR+ [11] ($\uparrow$), and LPIPS+ ($\downarrow$[34], respectively.
@@ 5.2 Ablation Study
We perform an ablation study to qualitatively and quantitatively validate the effect of threshold thrthr in Eq. (9) for generating mask ${\mathbf{M}}^{1}$. In Fig. 3, we show the mask propagation and editing results when thrthr is 5, 35, and 65, respectively. Here, $thr=5$ leads to a relatively large editing region in ${\mathbf{M}}^{1}$. Since our FREE-Edit does not perform feature injection in the editing region, the corresponding editing results of $thr=5$ still present unsatisfactory motion (see man’s hat). In contrast, $thr=65$ yields a smaller editing region in ${\mathbf{M}}^{1}$, which brings excessive feature injection from the source video, thus suffering from undesired content leakage. In contrast, $thr=35$ obtains the most accurate editing regions and the best editing results. Thus, we use $thr=35$ as the default setting. Besides, Tab. 1 reports CLIP score and Warp Error of the editing results of different thrthr. Here, $thr=5$ performs poorly in temporal coherence, and $thr=65$ is inferior in edit consistency since they correspond to larger and smaller editing masks, leading to insufficient and excessive injection.
5.3 Qualitative Results
In Fig. 4, we first present a qualitative comparison between FREE-Edit w/o injection, w/ vanilla injection, and w/ REE injection, where the first-frame mask is also listed. FREE-Edit w/o injection cannot maintain the video motion well (see movement of people). FREE-Edit w/ vanilla injection suffers from the issue of undesired content leakage. In contrast, our standard FREE-Edit can achieve the best performance, which demonstrates its capability of holding the non-editing area while performing effective editing propagation. In the last row of each example, we show the editing mask warped by the first-frame mask, which still corresponds to the editing area well throughout the entire video.
@@ 5.3 Qualitative Results
Then, in Fig. 5, we present a qualitative comparison with the current state-of-the-art image-driven video editing methods: Go-with-the-Flow [4], VideoShop [6], AnyV2V [12], and I2VEdit [19]. Note that VideoShop and I2VEdit can only process approximately 14 and 20 frames, respectively, on an NVIDIA A100 GPU. Following the official implementation of I2VEdit, we divide each video into segments of 14 and 20 frames for I2VEdit and VideoShop, respectively, where the last edited frame of the previous segment serves as the first edited frame for the current segment.
As shown in Fig. 5, I2VEdit [19] fails to maintain appearance consistency as the frame increases. The training-free methods, AnyV2V [12] and VideoShop [6], exhibit more severe degradation in visual quality. The very recent method Go-with-the-Flow [4] can present a good appearance, while facing inaccurate subtle motion (see human arm). In contrast, our FREE-Edit, with the proposed REE injection method, achieves the most stable and visually coherent results. Our approach effectively propagates the appearance from the first edited frame to subsequent frames. More examples are provided in our supplementary material.
Ours vs Editing Consistency $\uparrow$ Overall Quality $\uparrow$
AnyV2V 79.6% 81.9%
I2VEdit 78.5% 74.6%
VideoShop 85.2% 81.4%
Go-with-the-Flow 72.5% 68.4%
FREE-Edit w/o injection 82.2% 71.3%
FREE-Edit w/ vanilla injection 77.8% 69.8%

### 22 Mar 2026 v2

@@ Abstract
Image-driven video editing aims to propagate edit contents from the modified first frame to the remaining frames. Existing methods usually invert the source video to noise using a pre-trained image-to-video (I2V) model and then guide the sampling process using the edited first frame. Generally, a popular choice for maintaining motion and layout from the source video is intervening in the denoising process by injecting attention during reconstruction. However, such injection often leads to unsatisfactory results, where excessive injection leads to conflicting semantics with the source video while insufficient injection brings limited source representation. Recognizing this, we propose an Editing-awaRE (REE) injection method to modulate the injection intensity of each token. Specifically, we first compute the pixel difference between the source and edited first frame to form a corresponding editing mask. Next, we track the editing area throughout the entire video by using optical flow to warp the first-frame mask. Then, editing-aware feature injection intensity for each token is generated accordingly, where injection is not conducted in editing areas. Building upon REE injection, we further propose a zero-shot image-driven video editing framework with recent-emerging rectified-Flow models, dubbed FREE-Edit. Without fine-tuning or training, our FREE-Edit demonstrates effectiveness in various image-driven video editing scenarios, showing its capability to produce higher-quality outputs compared with existing techniques. Project page: https://free-edit.github.io/page/.
Figure 1: We propose an Editing-awaRE (REE) feature injection method for pre-trained rectified Flow models to conduct zero-shot image-driven video editing, dubbed FREE-Edit. Given the edited first frame, our method generates the output video, which propagates the editing contents while leaving other parts (e.g., motion) unchanged. The commonly used vanilla feature injection always leads to semantics conflicting with the edited first frame. We also provide the w/o injection results, which cannot preserve motion well. In contrast, our REE feature injection method adaptively modulates the injection intensity of each token, achieving the most elegant results.
@@ 1 Introduction
Given the above considerations, we propose an Editing-awaRE (REE) injection method, which modulates the injection intensity of each token according to whether it is edited or not. Specifically, we design a modulation weight $\bm{\lambda}$ to implement a linear interpolation with Query $\mathbf{Q}$} (or Key $\mathbf{K}$}) in the self-attention of the reconstruction branch and the corresponding $\tilde{{\mathbf{Q}}}$ (or $\tilde{{\mathbf{K}}}$) of the editing branch. To compute the modulation weight $\bm{\lambda}$, we first use the difference between the source and edited first frames to mask the editing area. Next, we track the editing area throughout the entire video by evaluating optical flow [31] for the source video, which is used to warp the first-frame mask. Then, the modulation weight $\bm{\lambda}$ can be obtained by inverting the warped mask sequence, where no injection is performed in the editing area to maintain appearance modification.
We apply the proposed REE injection method into the pre-trained Flow model LTX-Video [9] to form a training-free image-driven video editing framework, which inspires the name of our method: FREE-Edit. Though simple, the extensive evaluation of FREE-Edit demonstrates our capability of generating high-quality appearance- and temporally consistent outputs. For example, as seen in Fig. 1, motion cannot be preserved well in FREE-Edit w/o-injection due to the limited motion signal. Besides, FREE-Edit w/ vanilla feature injection always leads to conflicting semantics from the source video with the edited appearance. In contrast, the proposed REE injection yields the most visually consistent results. To sum up, our contributions are:
@@ 2 Related Works
Text-driven Video Generation and Editing. Early Text-to-Video (T2V) models  [1, 3] opt to insert temporal layers a pretrained Text-to-Image (T2I) model [27] to obtain temporal modeling. Another line of methods [9, 39] employs 3D VAE to incorporate 3D convolutions, which operate videos both spatially and temporally. Besides, text-driven video editing expects to generate an edited video that adheres to the target prompt while the original motion and semantic layout are preserved. Many efforts [37, 25, 7, 16, 18] expand the U-Net architecture of a pre-trained T2I diffusion model (e.g., SD [27]) along the temporal axis for 3D correspondence, where the denoising process is performed under the target-prompt guidance. Recent progress [33] has also included RF models to perform editing. Additionally, some methods use multiple frames [13, 35, 8, 38] or incorporate motion guidance [34, 29] for I2V generation.
Image-to-Video Generation and Editing. Several representative works [2, 9, 39] establish the foundation for image-to-video (I2V). Then, a group of methods [9, 32] propose a first-frame conditioning model with rectified-flow training procedure.Besides, some techniques [14, 6, 21] use these I2V models to carry out image-driven video editing by propagating the first edited frame. Here, besides training-free AnyV2V [14] and VideoShop [6], I2VEdit [21] performs finetuning on temporal attention layers for coarse motion extraction, which, however, is time- and resource-consuming. For example, I2VEdit takes about 2 hours for a 40-frame video in an NVIDIA A100 GPU, which is unfavourable in real applications. Moreover, very recent work Go-with-the-Flow [4] proposes a motion-controlled method by fine-tuning video diffusion models, which also allows for first-frame guided video editing.
@@ 3 Preliminary
After the velocity field $v_{\theta}$ is trained, pure noise ${\bm{z}}_{1}$ is gradually recovered as a clean latent ${\bm{z}}_{0}$ using NN discrete timesteps $t_{i}$:
${\bm{z}}_{t_{i-1}}={\bm{z}}_{t_{i}}+(t_{i-1}-t_{i})v_{\theta}({\bm{z}}_{t_{i}},t_{i}).$ (3)
Transformer Model for Image-to-Video. LTX-Video [9] is an open-source real-time generative model for I2V and T2V. Specifically, it uses a training procedure based on RF with 28 DiT transformer blocks [23]. Besides, it employs 3D causal VAEs to deal with spatio-temporal representation, where a $(1+rn)$-frame video is compressed into $(1+n)$ latent frames with temporal compression ratio $r$. Here, the first frame is encoded as a separate latent frame. Additionally, LTX-Video can perform real-time generation due to its highly compressed latent space.
Figure 2: The pipeline illustration of our FREE-Edit. Top: It obeys an “inversion-then-editing” pipeline. Starting from inverted noisy latent ${\bm{z}}_{1}$, the reconstructed video ${\bm{z}}_{0}$ and edited video $\tilde{{\bm{z}}}_{0}$ take the source ${{\mathbf{X}}}^{1}$ and edited first frame $\hat{{\mathbf{X}}}^{1}$ as conditions, respectively. Bottom: We design an Editing-awaRE (REE) injection method, which designs a modulation weight $\bm{\lambda}$ to adaptively replace the intermediate model representations ($\tilde{{\mathbf{Q}}}$ and $\tilde{{\mathbf{K}}}$) in the editing process with those ($\mathbf{Q}$} and $\mathbf{K}$}) in the reconstruction process through self-attention blocks. Here, we first use optical flow to warp the automatically calculated first-frame editing mask, which yields tracked editing masks for subsequent frames. Based on them, we compute the modulation weight $\bm{\lambda}$ for each token, where no injection is performed in editing regions.
@@ 4 Method
The rest of this section is organized as follows. In Sec. 4.1, we first present a baseline editing framework that employs the vanilla feature injection on a pre-trained video RF model (e.g., LTX-Video [9]). Then, we elaborate on the proposed Editing-awaRE (REE) feature injection method and our FREE-Edit framework in Sec. 4.2.
@@ 4.1 Vanilla Feature Injection on RF Models
Considering motion signals from the inverted noise alone is limited, AnyV2V[14] introduces a vanilla feature injection method in “inversion-then-editing” pipeline. Next, we will present how to reproduce it to match RF video models. Formally, we use $\mathbf{A}$} and $\tilde{{\mathbf{A}}}$ to denote the full spatiotemporal self-attention maps in source denoising (reconstruction) and editing denoising trajectory, respectively. Then, to obtain more layout and motion representation from the source video, vanilla injection applies the attention maps $\mathbf{A}$}} to replace attention maps $\tilde{{\mathbf{A}}}$, where $\tilde{{\mathbf{A}}}$ and $\mathbf{A}$} are obtained by:
@@ 4.2 The Proposed REE Injection and FREE-Edit
The remaining problem is how to use the editing mask sequence ${\mathbf{M}}=\{{\mathbf{M}}^{0},{\mathbf{M}}^{1},...,{\mathbf{M}}^{rn}\}$ to generate modulation weight $\bm{\lambda}$. Since the temporal compression of the RF I2V model, we take the mask of the $\left\lceil\frac{r}{2}\right\rceil$-th frame for each $r$-frame video chunk. Thus, the compressed editing mask sequence ${\mathbf{M}}_{\mathbf{c}}=\{{\mathbf{M}}^{0},{\mathbf{M}}^{\left\lceil\frac{r}{2}\right\rceil},...,{\mathbf{M}}^{\left\lceil\frac{r}{2}\right\rceil\cdot n}\}$ is with $(1+n)$ mask frames. Then, to match the dimension of Query and Key in Eq. (7), we further down-sample and flatten each mask frame in ${\mathbf{M}}_{\mathbf{c}}$ into size of $l\times 1$, yielding ${\mathbf{M}}_{c\cap d}\in\mathbb{R}^{(1+n)\times l\times 1}$. Finally, our modulation weight $\bm{\lambda}$ can be obtained by inverting the ${\mathbf{M}}_{c\cap d}$:
@@
FREE-Edit With Our REE Injection. In Fig. 2, we propose a training-free image-driven video editing framework, dubbed FREE-Edit, which obeys an “inversion-then-editing” pipeline. The core is our REE injection, which adaptively replaces $\tilde{{\mathbf{Q}}}$ and $\tilde{{\mathbf{K}}}$ in the editing trajectory with $\mathbf{Q}$}} and $\mathbf{K}$}} in the reconstruction editing trajectory through self-attention blocks in transformers. Here, clean tokens (i.e., source reference tokens and edited reference tokens corresponding to ${{\mathbf{X}}}^{1}$ and $\hat{{\mathbf{X}}}^{1}$) are combined with noisy tokens in each time step.
Note that text-driven technique Fatezero [25] also explores attention blending with masks, which are generated from cross-attention maps of the text token associated with the foreground object. However, cross-attention maps struggle to indicate editing regions in certain editing types (e.g., adding objects), where the location of the added object is uncertain. In contrast, we use optical flow to track editing regions, which can be well adapted to various editing types.
Our method also differs from TokenFlow [7], which propagates diffusion features during editing based on inter-frame correspondences calculated during inversion. Specifically, TokenFlow performs correspondence propagation in text-driven editing, whereas FREE-Edit employs editing-mask propagation in image-driven editing.
@@ 5.1 Experimental Setup
Implementation Details. We conduct all experiments on a single Tesla A100 GPU using PyTorch [22]. We use the classifier-free guidance (CFG) [11] of 3 during the editing process, and that of 1 in the inversion process, which is under the same negative prompt “worst quality, inconsistent motion, blurry, jittery, distorted”. Additionally, we adopt 50 time steps as noise schedule. We set $thr=35$ in Eq. (9) empirically. We use LTX-Video-2B [9] for the base I2V model for its real-time performance. As for feature injection, we conduct Eq. (7) in all transformer blocks and all time steps.
I2V-Edit-Benchmark. Considering there is no publicly available dataset for image-driven video editing, this paper constructs a mini benchmark named I2V-Edit-Bench to evaluate the proposed method. Specifically, it contains 60 videos and their corresponding edited first frames. The videos are collected from the Internet website, and DAVIS dataset [24], and their edit frames are created by using professional software (e.g., Photoshop) and MagicBrush [40]. Each video consists of 41 to 200 frames. We provide 6 editing types: including object replacement, style transfer, adding/removing objects, changing attributes, and background editing.
@@
Evaluation Metrics of Edited Videos. We use CLIP score ($\uparrow$) to measure editing fidelity, which calculates the similarity of CLIP image embedding [26] of each edited frame and the given first edit frame. Besides, we apply Warp Error ($\downarrow$[7, 16] to reflect temporal consistency. Specifically, we estimate optical flow [31] of the input video to warp the edited frames and compute average MSE between warped frames and the target ones. Moreover, with our generated mask sequence $\mathbf{M}$}, we measure the fidelity of the non-editing regions by five metrics: CLIP score+ ($\uparrow$), Warp Error+ ($\downarrow$), SSIM+ [36] ($\uparrow$), PSNR+ [12] ($\uparrow$), and LPIPS+ ($\downarrow$[41], respectively.
@@ 5.2 Ablation Study
We perform an ablation study to qualitatively and quantitatively validate the effect of threshold thrthr in Eq. (9) for generating mask ${\mathbf{M}}^{1}$. In Fig. 3, we show the mask propagation and editing results when thrthr is 5, 35, and 65, respectively. Here, $thr=5$ leads to a relatively large editing region in ${\mathbf{M}}^{1}$. Since our FREE-Edit does not perform feature injection in the editing region, the corresponding editing results of $thr=5$ still present unsatisfactory motion (see man’s hat). In contrast, $thr=65$ yields a smaller editing region in ${\mathbf{M}}^{1}$, which brings excessive feature injection from the source video, thus suffering from undesired content leakage. In comparison, $thr=35$ obtains the most accurate editing regions and the best editing results. Thus, we use $thr=35$ as the default setting. Besides, Tab. 1 reports CLIP score and Warp Error of the editing results of different thrthr. Here, $thr=5$ performs poorly in temporal coherence, and $thr=65$ is inferior in edit consistency since they correspond to larger and smaller editing masks, leading to insufficient and excessive injection.
@@
5.3 Qualitative Results
In Fig 4, we first present a qualitative comparison between FREE-Edit w/o injection, w/ vanilla injection, and w/ REE injection, where the first-frame mask is also listed. FREE-Edit w/o injection cannot maintain the video motion well (see movement of people). FREE-Edit w/ vanilla injection suffers from the issue of undesired content leakage. In contrast, our standard FREE-Edit can achieve the best performance, which demonstrates its capability of preserving the non-editing area while performing effective editing propagation. In the last row of each example, we show the editing mask warped by the first-frame mask, which still corresponds to the editing area well throughout the entire video.
Then, in Fig. 5, we present a qualitative comparison with state-of-the-art image-driven video editing methods: Go-with-the-Flow [4], VideoShop [6], AnyV2V [14], and I2VEdit [21]. Note that VideoShop and I2VEdit can only process around 14 and 20 frames respectively on an NVIDIA A100 GPU. Following the official implementation of I2VEdit, we split each video into chunks of 20 frames for VideoShop and 14 frames for I2VEdit, where the last edited frame of the previous chunk is used as the initialization for the current chunk.
As shown in Fig. 5, I2VEdit [21] fails to maintain appearance consistency as the frame increases. The training-free methods, AnyV2V [14] and VideoShop [6], exhibit more severe degradation in visual quality. The very recent method Go-with-the-Flow [4] can present a good appearance, while facing inaccurate subtle motion (see human arm). In contrast, our FREE-Edit, with the proposed REE injection method, yields the most stable and visually coherent results. Our approach effectively propagates the appearance from the first edited frame to subsequent frames. More examples are provided in the Appendix.
Ours vs Editing Consistency $\uparrow$ Overall Quality $\uparrow$
AnyV2V 79.6% 81.9%
I2VEdit 78.5% 74.6%
VideoShop 85.2% 81.4%
Go-with-the-Flow 72.5% 68.4%
FREE-Edit w/o injection 82.2% 71.3%
FREE-Edit w/ vanilla injection 77.8% 69.8%